OpenStudy (anonymous):

I GIVE MEDALS. Two asteroids are 75,000 m apart. One has a mass of 8 107 kg. If the force of gravity between them is 1.14 N, what is the mass of the other asteroid? A. 3.4 X 10^11 kg B. 1.2 X 10^12 kg C. 1.2 X 10^10 kg D. 8.3 X 10^12 kg

OpenStudy (anonymous):

Haven't we done this once? F = k m M / r^2 etc.

OpenStudy (cybershadow):

ok , We have been given r = distance between the asteroids = 75000 m and mass of one asteroid (m1) = 107 kg m2 = mass of second asteroid (to be found) and if the force of gravity between them ,is F = 1.14 N also we know that Force of gravity is F = Gm1m2/r^2 therefore, 1.14= G*107*m2/(75000)^2 m2 = 1.14 * 75000 * 75000 /107G (lemme calculate this :P, and btw i m guessing that it is a typo in your question u have type mass 8 107 kg <--- i m guessing that extra is a typo rite?) :P

OpenStudy (anonymous):

Its supposed to be 8X10^7 kg, sorry @CyberShadow

OpenStudy (cybershadow):

Gothca! Ok so that makes our equation m2 = 1.14 * (75000)^2 /6.6 * 10^-11 * 8 * 10 ^7 so after all the calculatioin m2 = 121.44 * 10^10 (approximately) i m not sure abt my calculation tho ,, u can try calculating it on ur own :)

OpenStudy (anonymous):

Is there any way to simplify it? Those are way off from my choices. Ill calculate it and see if i get the same thing tho @CyberShadow

OpenStudy (cybershadow):

actually the calculation is corect

OpenStudy (cybershadow):

121.4 * 10^10 = 1.2 * 10 ^12

OpenStudy (cybershadow):

sorry dint read ur options first time

OpenStudy (anonymous):

It's fine! how did you multiply 10^10? @CyberShadow

OpenStudy (cybershadow):

ok see , we had 121.4 *10 ^10 right? then in the options we have 1.2 so i converted 121.4 into 1.2 and how can we do that? by diving it by 100 i.e 10^2 and if we r diving it by 100 , we also have to multiply it by 100 so \[121.4\div(10^{2})\times10^{2}*10^{10}\] so that make it 1.2 * 10 ^ 12 got it?

OpenStudy (anonymous):

How did you get 10^2, was it just random? @CyberShadow

OpenStudy (cybershadow):

what wud u do, to convert 121.4 to 1.2 ?

OpenStudy (anonymous):

oh, you divide. alright i got it cx thanks! @CyberShadow

OpenStudy (cybershadow):

No probs :)